OpenStudy (anonymous):

There are six numbers on a normal die. What is the probability that if you roll the die, you’ll get a number less than 3? Student Answer: 0.333 0.5<-my answer 0.667 0.75

OpenStudy (mertsj):

There are 6 possible outcomes Two of them are less than 3 P(less then 3)=2/6=1/3=.3333...

OpenStudy (anonymous):

ur right, the answer is 0.5